OpenStudy (anonymous):

Which Answer is correct? where {A} 5/2*6-3^2 = .8333.... or {B} 5/2*6-3^2=8.5885 I solved {A} By using the natural flow of brakets working my way out from the inered most set (5/(2*(6-(3^2)))) and other using PEMDAS

OpenStudy (anonymous):

Any kind of input is greatly needed, the answers I have gotten are completely different. I asked my teacher and a mathematician and was tol two different ways of doing this.

Directrix (directrix):

I did not get either answer. 5/2 * 6 - 3^2 = (5/2 * 6) - 3^2 = 30/2 - 9 = 15 - 9 = 6.